OpenStudy (anonymous):

the ratio of reds to greens was 5 to 16 and six times the number of reds exceeded the number of greens by 112. how many were reds and how many were greens? thanks =)

OpenStudy (cwrw238):

r/g = 5/16 6r = g + 112

OpenStudy (cwrw238):

solve this sytem of equations

OpenStudy (cwrw238):

from first equation g = 16r/5 now plug this into second equation

OpenStudy (anonymous):

r=5?

OpenStudy (anonymous):

16(5)/5?

OpenStudy (cwrw238):